OpenStudy (anonymous):

what is the anti-derivative of xe^-x^2

OpenStudy (anonymous):

(-1/2)e^-x^2

myininaya (myininaya):

\[\int\limits_{}^{}xe^{-x^2} dx\] let \[u=-x^2 => du=-2x dx\] \[\int\limits_{}^{}xe^{-x^2} dx=\frac{-1}{2}\int\limits_{}^{}e^u du\]

OpenStudy (anonymous):

hmmm i smell a u-sub try \[u=-x^2\] \[du=-2xdx\] and go from there

OpenStudy (ybarrap):

d/dx(e^(-x^2) x) = e^(-x^2) (d/dx(x))+x (d/dx(e^(-x^2))) = x (e^(-x^2) (d/dx(-x^2)))+e^(-x^2) (d/dx(x)) = e^(-x^2) x (d/dx(-x^2))+1 e^(-x^2) = e^(-x^2) x (-(d/dx(x^2)))+e^(-x^2) = e^(-x^2)-e^(-x^2) x (2 x)